OpenStudy (anonymous):

a paper in a scholarly journal once claimed every year since 1950 the number of American children gunned down has doubled. in 1950 just one child was gunned down. make a table of the number of children killed in each of the next 10 years take the logarithm of your counts and plot these logaithms against the year. from your graph find the approximate values of the slope and the intercept for the line. use the equation y=a+bx to predict the logarithm of the count for the 45th year

OpenStudy (anonymous):

not that hard, but very macob

OpenStudy (anonymous):

look at it like this, the 1st year is 1, the next is 2, the thirst is 4, and the fourth is 5th, and so on

OpenStudy (anonymous):

you can see its rises exponentially, try doing the rest your self

OpenStudy (anonymous):

if I hadn't tried myself already I wouldn't have posted it here

OpenStudy (♪chibiterasu):

1 2 4 8 16 32 64 128 256 512 <<< answer

OpenStudy (anonymous):

and how is that the answer? it says for the 45th year

OpenStudy (anonymous):

can someone please help me instead of looking at it for 2 seconds. I'm so confused!!!!

OpenStudy (mathmale):

I'd model this exponential growth function by n=n0*e^kt, where n0 is the initial value (1) of the number of children gunned down, e, k is the growth constant, and t is the time (in years). Thus, n=e^kt. Using the fact that the number of such children doubles annually, we can find k by letting n=2 and t=1 (year): 2=e^k. Thus, n=e^kt becomes n=2^t. Taking the natural log of both sides, we get the log model y = ln n = ln 2^t =t*ln 2, or y = ln n = (ln 2)*t + 0, which has the form y = ax + b. Here the slope of the linear model is ln 2 and the y-intercept is 0 (zero). Check: if t = 1 (yr), ln n = ln 2, or n = 2 (correct); if t = 2, then ln n = (ln 2)*2 = ln 4, so that n = 4 (as expected). Thus, for the 45th year, ln n = (ln 2)*45, which implies n = 2*45. We were supposed to "predict the logarithm of the count," which is ln n = (ln 2)*45.

OpenStudy (anonymous):

If it were 2^45 that would mean 3.5 x 10^13 children gunned down in 45th year. Seems high! US population is 3 x 10^8.